### v1

# Non-perturbative self-consistent electron-phonon spectral functions and transport

Abstract
Electron-phonon coupling often dominates the electron spectral functions and carrier transport properties. However, studies of this effect in real materials have largely relied on perturbative one-shot methods due to the lack of a first-principles theoretical and computational framework. Here, we present a self-consistent theory and implementation for the non-perturbative calculations of spectral functions and conductivity due to electron-phonon coupling. Applying this method to monolayer InSe, we demonstrate that self-consistency qualitatively affects the spectral function and transport properties compared to state-of-the-art one-shot calculations and allow one to reconcile experimental angle-resolved photoemission experiments. The developed method can be widely applied to materials with dominant electron-phonon coupling at moderate computational cost.
The spectral function is a key quantity in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
Ab initio calculation of the electron-phonon self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17] and solve the Dyson equation [10] or apply the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25] to calculate the spectral functions. However, perturbative approaches often fail, leading to divergent quasiparticle dispersion [26, 27, 25] or unphysical spectral properties [25]. Importantly, this problem is not restricted to strongly coupled systems. The negative curvature in the quasiparticle dispersion of the Fröhlich model [26] and the divergent linewidth of piezoelectric materials [28] are examples where perturbation theories fail. More sophisticated methods that include higher-order electron-phonon coupling [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] are limited to model Hamiltonians. Non-perturbative methods based on molecular dynamics [44, 45] or special-displacement supercells [46, 47] are difficult to converge [45, 48], and cannot describe non-adiabatic electron-phonon coupling which is crucial in polar materials [27, 49].
The electron spectral function also describes the transport of charged carriers. Using the Green–Kubo (GK) theory [50, 51], one can compute the conductivity without current vertex correction directly from the spectral functions [52, 53], which is sometimes called the Kubo formula in the bubble approximation or the independent-particle approximation [54]. This formulation has the advantage of including broadening and satellites of spectral functions beyond the quasiparticle approximation [55, 56, 57, 58, 45].
A material of particular interest for the non-perturbative electron-phonon coupling is monolayer InSe. While monolayer InSe is known for its high electron mobility [59, 60], its valence band has a flat, Mexican-hat-like dispersion [61, 62, 63, 64, 65, 66, 67], which leads to strong electron-phonon interaction. In particular, state-of-the-art one-shot calculations of the spectral function yield a dynamical splitting of a single valence band into two bands with a gap larger than 150 meV [68]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [69, 70, 71] and the GK method with the one-shot FM spectral function [68]. Since the renormalization is large, non-perturbative corrections are expected to play an important role in InSe.
In this work, we present a non-perturbative ab initio calculation of the electron-phonon spectral function and conductivity. We use the self-consistent Fan–Migdal (scFM) method [35, 36, 37, 38, 39, 40, 41, 42, 43], which self-consistently updates the self-energy with the renormalized Green’s function. Using the self-consistent spectral functions, we then use the GK method to compute the conductivity. We find that self-consistency significantly affects the spectral function, direct current (DC) conductivity, and optical conductivity of monolayer InSe. Such an effect is present in any materials and can be studied using this scFM method.
The most used approximation for the electron-phonon self-energy is the FM formula [2]:
where is the volume of the Brillouin zone, the electron-phonon matrix element, and the electron and phonon energies, the Fermi–Diarc distribution function, , and the Bose-Einstein distribution function. We neglect the band-off-diagonal self-energy, which is negligible if the bands do not hybridize [74], as studied here. In this work, we assume that the bare electron Green’s function is obtained from Kohn–Sham density functional theory and that the phonon Green’s function and electron-phonon coupling are obtained from adiabatic density functional perturbation theory [75, 76, 2]. We neglect the renormalization of the phonon Green’s function.
Using the FM self-energy, the renormalized electron dispersion can be computed at the level of Rayleigh–Schrödinger (RS) perturbation theory [77, 78, 27, 2, 23] by adding the on-shell FM self-energy to the bare band energy:
The energy-dependent spectral function can be calculated by solving the Dyson equation:
To capture higher-order electron-phonon coupling, one often uses the cumulant approximation, which replaces the Dyson equation with the cumulant expansion
where is the Heaviside function. One then approximates the cumulant function using second-order perturbation theory [21]:
Unlike the FM approximation, which gives a single satellite [15], the cumulant approximation can generate multiple satellite peaks as observed in experiments [11, 12, 13, 14]. However, the cumulant approximation can lead to unphysically large renormalizations for -points away from the band edge [25, 43], as does the RS [26] approximation.
To include the non-perturbative effects of electron-phonon coupling, we go one step further and use the self-consistent Fan–Migdal (scFM) method. The scFM self-energy is obtained by replacing the bare electron Green’s function in the FM approximation with the dressed one. The imaginary part of the scFM self-energy is [57]:
The real part is then obtained from the Kramers–Kronig relation:
where denotes the Cauchy principal value. The spectral function in Eq. (7) is obtained by solving the Dyson equation:
A self-consistent solution of Eqs. (7-10) gives the scFM self-energy and spectral function. For metals or doped semiconductors, we iteratively update the temperature-dependent chemical potential to match the total electron occupation.
We remark that it is advantageous to calculate the imaginary part first and then compute the real part using the Kramers–Kronig relation. In fact, the relation between the real parts of and involves a -dependent energy integral, see Eq. (S46) of the supplemental of Ref. [28]. The separation of Eq. (7) and Eq. (8) allows to perform the -integration for the imaginary part first, and then perform a -independent energy integral after.
In practical ab initio calculations, several hundred bands per atom are needed to converge the real part of the FM self-energy [48]. Calculating such a term via the Kramers–Kronig transformation is impractical since it requires calculating up to eV. We simplify the calculation by exploiting the fact that the contribution of high-energy bands to the self-energy at small can be viewed as a static correction, which can be computed via the self-consistent solution of linear Sternheimer equations [77, 78]. We self-consistently compute considering only the states inside the active energy window . We approximate the remainder to be static and compute it using Wannier function perturbation theory (WFPT) [79]:
Here, is the FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue , and is the Debye–Waller self-energy [5, 4, 3].
In practice, we initialize the scFM self-energy with from Eq. (1) with meV. We iterate until the maximum change of the self-energy for all states and energies (with respect to the band edge) is below meV. We have implemented scFM as an extension of our ElectronPhonon.jl package [80], which relies on quantities computed using Quantum ESPRESSO [81], Wannier90 [82], and EPW [83, 84]. Computational details are provided in Sec. S1 of the SM [72].
The scFM method, also known as the self-consistent Born approximation or the noncrossing approximation, has been applied to the Holstein [35, 36, 37, 39, 40, 41, 42, 43] and Anderson–Holstein models [38]. State-of-the-art ab initio calculations [57, 58] used a one-shot calculation of Eqs. (7) and (8) to perturbatively compute the electron-phonon self-energy using the spectral function obtained from dynamical mean-field theory calculations. Here, we self-consistently solve Eqs. (7-10) to include non-perturbative effects of electron-phonon coupling on the spectral function in real materials.
We apply the scFM method to monolayer InSe which has a Mexican-hat-like valance band [61, 62, 63, 65]. Due to this dispersion, hole-doped InSe monolayers exhibit strong electron-phonon effects. In particular, the one-shot FM spectral functions show a dynamic splitting of a single valence band into two peaks with a gap of 150 meV or larger [68]. We study whether this large renormalization persists under a self-consistent update of the spectral function.
Figures 1(a-d) show the valence band spectral functions of monolayer InSe at 100 K which is compared to the monolayer angle-resolved photoemission spectra (ARPES) intensity from Ref. [73] measured at the same temperature. We first find that the RS dispersion has an inverted curvature at the VBM, an artifact also observed for the Fröhlich model with strong electron-phonon coupling [26]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot FM method yields two dynamically split bands with a splitting of 200 meV at the VBM, consistent with Ref. [68]. However, this large splitting vanishes under self-consistency. The scFM spectral function shows a broadened quasiparticle peak with multiple satellites. The distance between the satellite peaks ranges from 30 to 50 meV and is comparable to the LO phonon energy of 27 meV. An in-depth analysis of the difference between FM and scFM for different temperatures is provided in Sec. S2 of the SM [72]. Our scFM results are consistent with the absence of a 150 meV scale gap [65, 73] and the flatness of the band shown in Fig.1(e), in better agreement than the cumulant result. We expect that higher-resolution ARPES will observe the satellites predicted by our scFM calculation.
Beyond ARPES predictions that can sometimes be challenging to verify experimentally, the electron spectral function can be directly used to compute carrier conductivity, which is a much simpler quantity to measure. Using the GK linear response theory and neglecting the current vertex correction, the conductivity is [52]
The corresponding mobility is with the carrier concentration. Equation (12) has been used to calculate the conductivity due to electron-electron interaction [53], electron-phonon interaction [55, 56, 45], or both [57, 58]. Here, we study how the self-consistent renormalization of the spectral function affects the DC and optical conductivity.
Figure 2(a) shows the intrinsic DC mobility of InSe. While the Boltzmann transport equation (BTE) and self-energy relaxation time approximation (SERTA) mobilities [85] are almost the same and agree with those of Ref. [70, 71], the GK mobility with the scFM spectral function is 3–4 times larger. This result implies that the quasiparticle approximation leads to a significant underestimation of the mobility. We also find that using the one-shot FM spectral function instead of the scFM leads to an overestimation by more than an order of magnitude. These results can be understood in terms of the effective electron lifetimes shown in Fig. 2(b). The SERTA and BTE mobilities are linked to the inverse electron lifetimes, which are proportional to the imaginary part of the FM self-energy at the bare band energy,
For the GK mobility, we define the effective inverse lifetimes using the self-energy at the quasiparticle energy instead of the bare band energy:
Here, the quasiparticle energy and are the energy at which the corresponding spectral functions [Eq. (3)] and [Eq. (9)] are maximal. Near the VBM, the BTE inverse lifetimes have a peak, while the scFM ones show a smoother variation, due to self-consistent broadening. The lower scattering rates increase the mobility. For the one-shot FM self-energy, the scattering rates near the VBM are tiny, because the renormalized valence bands lie within the DFT band gap where the one-shot FM self-energy has a negligible imaginary part. This is an artifact of the one-shot calculation.
Finally, we show in Fig. 3 the alternating current (AC) mobility using the FM and scFM self-energy and the GK method. For the one-shot FM spectral function, we find a peak around 200–300 meV, corresponding to the gap shown in Fig. 1(c) [68]. This feature is found to be stable regardless of the broadening parameter (see Fig. LABEL:supp-fig:ac_sc_broadening of the SM [72]). However, when the scFM spectral function is used, the unphysical peaks are washed out due to the reduction of the splitting energy and the broadening of the Drude peak.
The self-consistent framework developed in this work provides a versatile tool for analyzing the effect of higher-order electron-phonon coupling with moderate computational effort, filling the gap between single-shot perturbative approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, our framework can be extended to other properties such as superconductivity and indirect absorption [86, 87]. Our work lays the groundwork for applying diagrammatic many-body approaches to electron-phonon systems. The scFM method could be extended to study contributions beyond the Fan–Migdal diagram, such as the dynamical screening of phonons [35, 36, 88, 89, 90, 91] and the electron-phonon vertex correction [39, 41, 92] in real materials. Combining the scFM formalism with dynamical mean-field theory [93, 94] and its diagrammatic extensions [95] to study the interplay of electron-electron [53] and electron-phonon coupling [96] on transport beyond the one-shot approximation [57, 58] is another promising direction for future research.
Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73]. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247.References
- Damascelli et al. [2003] A. Damascelli, Z. Hussain, and Z.-X. Shen, Angle-resolved photoemission studies of the cuprate superconductors, Rev. Mod. Phys. 75, 473 (2003).
- Giustino [2017] F. Giustino, Electron-phonon interactions from first principles, Rev. Mod. Phys. 89, 015003 (2017).
- Allen and Cardona [1983] P. B. Allen and M. Cardona, Temperature dependence of the direct gap of Si and Ge, Phys. Rev. B 27, 4760 (1983).
- Allen and Cardona [1981] P. B. Allen and M. Cardona, Theory of the temperature dependence of the direct gap of germanium, Phys. Rev. B 23, 1495 (1981).
- Allen and Heine [1976] P. B. Allen and V. Heine, Theory of the temperature dependence of electronic band structures, J. Phys. Condens. Matter 9, 2305 (1976).
- Marini [2008] A. Marini, Ab initio finite-temperature excitons, Phys. Rev. Lett. 101, 106405 (2008).
- Giustino et al. [2010] F. Giustino, S. G. Louie, and M. L. Cohen, Electron-phonon renormalization of the direct band gap of diamond, Phys. Rev. Lett. 105, 265501 (2010).
- Valla et al. [1999] T. Valla, A. V. Fedorov, P. D. Johnson, and S. L. Hulbert, Many-body effects in angle-resolved photoemission: Quasiparticle energy and lifetime of a Mo(110) surface state, Phys. Rev. Lett. 83, 2085 (1999).
- Eiguren et al. [2003] A. Eiguren, S. de Gironcoli, E. V. Chulkov, P. M. Echenique, and E. Tosatti, Electron-phonon interaction at the Be(0001) surface, Phys. Rev. Lett. 91, 166803 (2003).
- Engelsberg and Schrieffer [1963] S. Engelsberg and J. R. Schrieffer, Coupled electron-phonon system, Phys. Rev. 131, 993 (1963).
- Moser et al. [2013] S. Moser, L. Moreschini, J. Jaćimović, O. S. Barišić, H. Berger, A. Magrez, Y. J. Chang, K. S. Kim, A. Bostwick, E. Rotenberg, L. Forró, and M. Grioni, Tunable polaronic conduction in anatase , Phys. Rev. Lett. 110, 196403 (2013).
- Chen et al. [2015] C. Chen, J. Avila, E. Frantzeskakis, A. Levy, and M. C. Asensio, Observation of a two-dimensional liquid of Fröhlich polarons at the bare SrTiO3 surface, Nat. Commun. 6, 8585 (2015).
- Cancellieri et al. [2016] C. Cancellieri, A. S. Mishchenko, U. Aschauer, A. Filippetti, C. Faber, O. Barišić, V. Rogalev, T. Schmitt, N. Nagaosa, and V. N. Strocov, Polaronic metal state at the LaAlO3/SrTiO3 interface, Nat. Commun. 7, 10386 (2016).
- Wang et al. [2016] Z. Wang, S. McKeown Walker, A. Tamai, Y. Wang, Z. Ristic, F. Y. Bruno, A. De La Torre, S. Riccò, N. C. Plumb, M. Shi, P. Hlawenka, J. Sánchez-Barriga, A. Varykhalov, T. K. Kim, M. Hoesch, P. D. C. King, W. Meevasana, U. Diebold, J. Mesot, B. Moritz, T. P. Devereaux, M. Radovic, and F. Baumberger, Tailoring the nature and strength of electron-phonon interactions in the SrTiO3(001) 2D electron liquid, Nat. Mater. 15, 835 (2016).
- Verdi et al. [2017] C. Verdi, F. Caruso, and F. Giustino, Origin of the crossover from polarons to Fermi liquids in transition metal oxides, Nat. Commun. 8, 15769 (2017).
- Migdal [1958] A. Migdal, Interaction between electrons and lattice vibrations in a normal metal, Sov. Phys. JETP 7, 996 (1958).
- Fan [1951] H. Y. Fan, Temperature dependence of the energy gap in semiconductors, Phys. Rev. 82, 900 (1951).
- Gunnarsson et al. [1994] O. Gunnarsson, V. Meden, and K. Schönhammer, Corrections to Migdal’s theorem for spectral functions: A cumulant treatment of the time-dependent Green’s function, Phys. Rev. B 50, 10462 (1994).
- Aryasetiawan et al. [1996] F. Aryasetiawan, L. Hedin, and K. Karlsson, Multiple plasmon satellites in Na and Al spectral functions from ab initio cumulant expansion, Phys. Rev. Lett. 77, 2268 (1996).
- Guzzo et al. [2011] M. Guzzo, G. Lani, F. Sottile, P. Romaniello, M. Gatti, J. J. Kas, J. J. Rehr, M. G. Silly, F. Sirotti, and L. Reining, Valence electron photoemission spectrum of semiconductors: Ab initio description of multiple satellites, Phys. Rev. Lett. 107, 166401 (2011).
- Story et al. [2014] S. M. Story, J. J. Kas, F. D. Vila, M. J. Verstraete, and J. J. Rehr, Cumulant expansion for phonon contributions to the electron spectral function, Phys. Rev. B 90, 195135 (2014).
- Gumhalter et al. [2016] B. Gumhalter, V. Kovač, F. Caruso, H. Lambert, and F. Giustino, On the combined use of approximation and cumulant expansion in the calculations of quasiparticle spectra: The paradigm of Si valence bands, Phys. Rev. B 94, 035103 (2016).
- Nery et al. [2018] J. P. Nery, P. B. Allen, G. Antonius, L. Reining, A. Miglio, and X. Gonze, Quasiparticles and phonon satellites in spectral functions of semiconductors and insulators: Cumulants applied to the full first-principles theory and the Fröhlich polaron, Phys. Rev. B 97, 115145 (2018).
- de Abreu et al. [2022] J. C. de Abreu, J. P. Nery, M. Giantomassi, X. Gonze, and M. J. Verstraete, Spectroscopic signatures of nonpolarons: the case of diamond, Phys. Chem. Chem. Phys. 24, 12580 (2022).
- Kandolf et al. [2022] N. Kandolf, C. Verdi, and F. Giustino, Many-body Green’s function approaches to the doped Fröhlich solid: Exact solutions and anomalous mass enhancement, Phys. Rev. B 105, 085148 (2022).
- Larsen [1966] D. M. Larsen, Polaron energy spectrum, Phys. Rev. 144, 697 (1966).
- Poncé et al. [2015] S. Poncé, Y. Gillet, J. Laflamme Janssen, A. Marini, M. Verstraete, and X. Gonze, Temperature dependence of the electronic structure of semiconductors and insulators, J. Chem. Phys. 143, 102813 (2015).
- Lihm et al. [2024] J.-M. Lihm, S. Poncé, and C.-H. Park, Self-consistent electron lifetimes for electron-phonon scattering, Phys. Rev. B 110, L121106 (2024).
- Prokof’ev and Svistunov [1998] N. V. Prokof’ev and B. V. Svistunov, Polaron problem by diagrammatic quantum Monte Carlo, Phys. Rev. Lett. 81, 2514 (1998).
- Jeckelmann and White [1998] E. Jeckelmann and S. R. White, Density-matrix renormalization-group study of the polaron problem in the Holstein model, Phys. Rev. B 57, 6376 (1998).
- Berciu [2006] M. Berciu, Green’s function of a dressed particle, Phys. Rev. Lett. 97, 036402 (2006).
- Bonča et al. [2019] J. Bonča, S. A. Trugman, and M. Berciu, Spectral function of the Holstein polaron at finite temperature, Phys. Rev. B 100, 094307 (2019).
- Carbone et al. [2021] M. R. Carbone, D. R. Reichman, and J. Sous, Numerically exact generalized Green’s function cluster expansions for electron-phonon problems, Phys. Rev. B 104, 035106 (2021).
- Mahajan et al. [2024] A. Mahajan, P. J. Robinson, J. Lee, and D. R. Reichman, Structure and dynamics of electron-phonon coupled systems using neural quantum states, arXiv:2405.08701 (2024).
- Säkkinen et al. [2015a] N. Säkkinen, Y. Peng, H. Appel, and R. Van Leeuwen, Many-body Green’s function theory for electron-phonon interactions: Ground state properties of the Holstein dimer, J. Chem. Phys. 143, 234101 (2015a).
- Säkkinen et al. [2015b] N. Säkkinen, Y. Peng, H. Appel, and R. Van Leeuwen, Many-body Green’s function theory for electron-phonon interactions: The Kadanoff-Baym approach to spectral properties of the Holstein dimer, J. Chem. Phys. 143, 234102 (2015b).
- Rademaker et al. [2016] L. Rademaker, Y. Wang, T. Berlijn, and S. Johnston, Enhanced superconductivity due to forward scattering in fese thin films on SrTiO3 substrates, New J. Phys. 18, 022001 (2016).
- Chen et al. [2016] H.-T. Chen, G. Cohen, A. J. Millis, and D. R. Reichman, Anderson-Holstein model in two flavors of the noncrossing approximation, Phys. Rev. B 93, 174309 (2016).
- Esterlis et al. [2018] I. Esterlis, B. Nosarzewski, E. W. Huang, B. Moritz, T. P. Devereaux, D. J. Scalapino, and S. A. Kivelson, Breakdown of the Migdal-Eliashberg theory: A determinant quantum Monte Carlo study, Phys. Rev. B 97, 140501 (2018).
- Dee et al. [2019] P. M. Dee, K. Nakatsukasa, Y. Wang, and S. Johnston, Temperature-filling phase diagram of the two-dimensional Holstein model in the thermodynamic limit by self-consistent Migdal approximation, Phys. Rev. B 99, 024514 (2019).
- Dee et al. [2020] P. M. Dee, J. Coulter, K. G. Kleiner, and S. Johnston, Relative importance of nonlinear electron-phonon coupling and vertex corrections in the Holstein model, Commun. Phys. 3, 145 (2020).
- Mitrić et al. [2022] P. Mitrić, V. Janković, N. Vukmirović, and D. Tanasković, Spectral functions of the Holstein polaron: Exact and approximate solutions, Phys. Rev. Lett. 129, 096401 (2022).
- Mitrić et al. [2023] P. Mitrić, V. Janković, N. Vukmirović, and D. Tanasković, Cumulant expansion in the Holstein model: Spectral functions and mobility, Phys. Rev. B 107, 125165 (2023).
- Zacharias et al. [2020] M. Zacharias, M. Scheffler, and C. Carbogno, Fully anharmonic nonperturbative theory of vibronically renormalized electronic band structures, Phys. Rev. B 102, 045126 (2020).
- Quan et al. [2024] J. Quan, C. Carbogno, and M. Scheffler, Carrier mobility of strongly anharmonic materials from first principles, Phys. Rev. B 110, 235202 (2024).
- Zacharias and Giustino [2016] M. Zacharias and F. Giustino, One-shot calculation of temperature-dependent optical spectra and phonon-induced band-gap renormalization, Phys. Rev. B 94, 075125 (2016).
- Zacharias and Giustino [2020] M. Zacharias and F. Giustino, Theory of the special displacement method for electronic structure calculations at finite temperature, Phys. Rev. Res. 2, 013357 (2020).
- Poncé et al. [2024] S. Poncé, J.-M. Lihm, and C.-H. Park, Zero-point renormalization of the bandgap, mass enhancement, and spectral functions: Validation of methods and verification of first-principles codes, arXiv:2410.14319 (2024).
- Miglio et al. [2020] A. Miglio, V. Brousseau-Couture, E. Godbout, G. Antonius, Y.-H. Chan, S. G. Louie, M. Côté, M. Giantomassi, and X. Gonze, Predominance of non-adiabatic effects in zero-point renormalization of the electronic band gap, npj Comput. Mater. 6, 167 (2020).
- Green [1954] M. S. Green, Markoff random processes and the statistical mechanics of time-dependent phenomena. II. irreversible processes in fluids, J. Chem. Phys. 22, 398 (1954).
- Kubo [1957] R. Kubo, Statistical-mechanical theory of irreversible processes. I. general theory and simple applications to magnetic and conduction problems, J. Phys. Soc. Jpn. 12, 570 (1957).
- Mahan [2000] G. D. Mahan, Many-particle physics, 3rd ed. (Kluwer Academic / Plenum Publishers, New York, 2000).
- Basov et al. [2011] D. N. Basov, R. D. Averitt, D. van der Marel, M. Dressel, and K. Haule, Electrodynamics of correlated electron materials, Rev. Mod. Phys. 83, 471 (2011).
- Poncé et al. [2020] S. Poncé, W. Li, S. Reichardt, and F. Giustino, First-principles calculations of charge carrier mobility and conductivity in bulk semiconductors and two-dimensional materials, Rep. Prog. Phys. 83, 036501 (2020).
- Zhou and Bernardi [2019] J.-J. Zhou and M. Bernardi, Predicting charge transport in the presence of polarons: The beyond-quasiparticle regime in SrTiO3, Phys. Rev. Res. 1, 033138 (2019).
- Chang et al. [2022] B. K. Chang, J.-J. Zhou, N.-E. Lee, and M. Bernardi, Intermediate polaronic charge transport in organic crystals from a many-body first-principles approach, npj Comput. Mater. 8, 63 (2022).
- Abramovitch et al. [2023] D. J. Abramovitch, J.-J. Zhou, J. Mravlje, A. Georges, and M. Bernardi, Combining electron-phonon and dynamical mean-field theory calculations of correlated materials: Transport in the correlated metal , Phys. Rev. Mater. 7, 093801 (2023).
- Abramovitch et al. [2024] D. J. Abramovitch, J. Mravlje, J.-J. Zhou, A. Georges, and M. Bernardi, Respective roles of electron-phonon and electron-electron interactions in the transport and quasiparticle properties of , Phys. Rev. Lett. 133, 186501 (2024).
- Bandurin et al. [2017] D. A. Bandurin, A. V. Tyurnina, G. L. Yu, A. Mishchenko, V. Zólyomi, S. V. Morozov, R. K. Kumar, R. V. Gorbachev, Z. R. Kudrynskyi, S. Pezzini, et al., High electron mobility, quantum Hall effect and anomalous optical response in atomically thin InSe, Nat. Nanotechnol. 12, 223 (2017).
- Wan et al. [2023] W. Wan, R. Guo, Y. Ge, and Y. Liu, Carrier and phonon transport in 2d InSe and its Janus structures, J. Condens. Matter Phys. 35, 133001 (2023).
- Zólyomi et al. [2014] V. Zólyomi, N. D. Drummond, and V. I. Fal’ko, Electrons and phonons in single layers of hexagonal indium chalcogenides from ab initio calculations, Phys. Rev. B 89, 205416 (2014).
- Rybkovskiy et al. [2014] D. V. Rybkovskiy, A. V. Osadchy, and E. D. Obraztsova, Transition from parabolic to ring-shaped valence band maximum in few-layer GaS, GaSe, and InSe, Phys. Rev. B 90, 235302 (2014).
- Debbichi et al. [2015] L. Debbichi, O. Eriksson, and S. Lebégue, Two-dimensional indium selenides compounds: An ab initio study, J. Phys. Chem. Lett. 6, 3098 (2015).
- Magorrian et al. [2016] S. J. Magorrian, V. Zólyomi, and V. I. Fal’ko, Electronic and optical properties of two-dimensional InSe from a DFT-parametrized tight-binding model, Phys. Rev. B 94, 245431 (2016).
- Kibirev et al. [2018] I. A. Kibirev, A. V. Matetskiy, A. V. Zotov, and A. A. Saranin, Thickness-dependent transition of the valence band shape from parabolic to Mexican-hat-like in the MBE grown InSe ultrathin films, Appl. Phys. Lett. 112, 191602 (2018).
- Graham et al. [2020] A. J. Graham, J. Zultak, M. J. Hamer, V. Zolyomi, S. Magorrian, A. Barinov, V. Kandyba, A. Giampietri, A. Locatelli, F. Genuzio, N. C. Teutsch, C. Salazar, N. D. M. Hine, V. I. Fal’ko, R. V. Gorbachev, and N. R. Wilson, Ghost anti-crossings caused by interlayer umklapp hybridization of bands in 2D heterostructures, 2D Materials 8, 015016 (2020).
- Stepanov et al. [2022] E. A. Stepanov, V. Harkov, M. Rösner, A. I. Lichtenstein, M. I. Katsnelson, and A. N. Rudenko, Coexisting charge density wave and ferromagnetic instabilities in monolayer InSe, npj Comput. Mater. 8, 118 (2022).
- Lugovskoi et al. [2019] A. V. Lugovskoi, M. I. Katsnelson, and A. N. Rudenko, Strong electron-phonon coupling and its influence on the transport and optical properties of hole-doped single-layer InSe, Phys. Rev. Lett. 123, 176401 (2019).
- Das et al. [2019] P. Das, D. Wickramaratne, B. Debnath, G. Yin, and R. K. Lake, Charged impurity scattering in two-dimensional materials with ring-shaped valence bands: GaS, GaSe, InS, and InSe, Phys. Rev. B 99, 085409 (2019).
- Li et al. [2019] W. Li, S. Poncé, and F. Giustino, Dimensional crossover in the carrier mobility of two-dimensional semiconductors: The case of InSe, Nano Lett. 19, 1774 (2019).
- Poncé et al. [2023a] S. Poncé, M. Royo, M. Stengel, N. Marzari, and M. Gibertini, Long-range electrostatic contribution to electron-phonon couplings and mobilities of two-dimensional and bulk materials, Phys. Rev. B 107, 155424 (2023a).
- Note [1] See Supplemental Material [URL will be inserted by publisher], which includes Refs. [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114], for the computational details and additional results.
- Hamer et al. [2019] M. J. Hamer, J. Zultak, A. V. Tyurnina, V. Zólyomi, D. Terry, A. Barinov, A. Garner, J. Donoghue, A. P. Rooney, V. Kandyba, A. Giampietri, A. Graham, N. Teutsch, X. Xia, M. Koperski, S. J. Haigh, V. I. Fal’ko, R. V. Gorbachev, and N. R. Wilson, Indirect to direct gap crossover in two-dimensional inse revealed by angle-resolved photoemission spectroscopy, ACS Nano 13, 2136 (2019).
- Lihm and Park [2020] J.-M. Lihm and C.-H. Park, Phonon-induced renormalization of electron wave functions, Phys. Rev. B 101, 121102 (2020).
- Gonze and Lee [1997] X. Gonze and C. Lee, Dynamical matrices, born effective charges, dielectric permittivity tensors, and interatomic force constants from density-functional perturbation theory, Phys. Rev. B 55, 10355 (1997).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Phonons and related crystal properties from density-functional perturbation theory, Rev. Mod. Phys. 73, 515 (2001).
- Gonze et al. [2011] X. Gonze, P. Boulanger, and M. Côté, Theoretical approaches to the temperature and zero-point motion effects on the electronic band structure, Ann. Phys. 523, 168 (2011).
- Poncé et al. [2014] S. Poncé, G. Antonius, P. Boulanger, E. Cannuccia, A. Marini, M. Côté, and X. Gonze, Verification of first-principles codes: Comparison of total energies, phonon frequencies, electron-phonon coupling and zero-point motion correction to the gap between abinit and qe/yambo, Comput. Mater. Sci. 83, 341 (2014).
- Lihm and Park [2021] J.-M. Lihm and C.-H. Park, Wannier function perturbation theory: Localized representation and interpolation of wave function perturbation, Phys. Rev. X 11, 041053 (2021).
- Lihm [2024] J.-M. Lihm, ElectronPhonon.jl (2024).
- Giannozzi et al. [2017] P. Giannozzi et al., Advanced capabilities for materials modelling with Quantum ESPRESSO, J. Condens. Matter Phys. 29, 465901 (2017).
- Pizzi et al. [2020] G. Pizzi et al., Wannier90 as a community code: New features and applications, J. Condens. Matter Phys. 32, 165902 (2020).
- Poncé et al. [2016] S. Poncé, E. Margine, C. Verdi, and F. Giustino, EPW: Electron-phonon coupling, transport and superconducting properties using maximally localized Wannier functions, Comput. Phys. Commun. 209, 116 (2016).
- Lee et al. [2023] H. Lee, S. Poncé, K. Bushick, S. Hajinazar, J. Lafuente-Bartolome, J. Leveillee, C. Lian, J.-M. Lihm, F. Macheda, H. Mori, H. Paudyal, W. H. Sio, S. Tiwari, M. Zacharias, X. Zhang, N. Bonini, E. Kioupakis, E. R. Margine, and F. Giustino, Electron–phonon physics from first principles using the EPW code, npj Comput. Mater. 9, 156 (2023).
- Poncé et al. [2018] S. Poncé, E. R. Margine, and F. Giustino, Towards predictive many-body calculations of phonon-limited carrier mobilities in semiconductors, Phys. Rev. B 97, 121201 (2018).
- Noffsinger et al. [2012] J. Noffsinger, E. Kioupakis, C. G. Van de Walle, S. G. Louie, and M. L. Cohen, Phonon-assisted optical absorption in silicon from first principles, Phys. Rev. Lett. 108, 167402 (2012).
- Tiwari et al. [2024] S. Tiwari, E. Kioupakis, J. Menendez, and F. Giustino, Unified theory of optical absorption and luminescence including both direct and phonon-assisted processes, Phys. Rev. B 109, 195127 (2024).
- Berges et al. [2023] J. Berges, N. Girotto, T. Wehling, N. Marzari, and S. Poncé, Phonon self-energy corrections: To screen, or not to screen, Phys. Rev. X 13, 041009 (2023).
- Macheda et al. [2024] F. Macheda, F. Mauri, and T. Sohier, Ab initio Van der Waals electrodynamics: Polaritons and electron scattering from plasmons and phonons in BN-capped graphene, Phys. Rev. B 110, 115407 (2024).
- Lihm and Park [2024] J.-M. Lihm and C.-H. Park, Plasmon-phonon hybridization in doped semiconductors from first principles, Phys. Rev. Lett. 133, 116402 (2024).
- Krsnik et al. [2024] J. Krsnik, D. Novko, and O. S. Barišić, Superconductivity in two-dimensional systems enhanced by nonadiabatic phonon-production effects, Phys. Rev. B 110, L180505 (2024).
- Lee et al. [2020] N.-E. Lee, J.-J. Zhou, H.-Y. Chen, and M. Bernardi, Ab initio electron-two-phonon scattering in GaAs from next-to-leading order perturbation theory, Nat. Commun. 11, 1607 (2020).
- Georges et al. [1996] A. Georges, G. Kotliar, W. Krauth, and M. J. Rozenberg, Dynamical mean-field theory of strongly correlated fermion systems and the limit of infinite dimensions, Rev. Mod. Phys. 68, 13 (1996).
- Kotliar et al. [2006] G. Kotliar, S. Y. Savrasov, K. Haule, V. S. Oudovenko, O. Parcollet, and C. A. Marianetti, Electronic structure calculations with dynamical mean-field theory, Rev. Mod. Phys. 78, 865 (2006).
- Rohringer et al. [2018] G. Rohringer, H. Hafermann, A. Toschi, A. A. Katanin, A. E. Antipov, M. I. Katsnelson, A. I. Lichtenstein, A. N. Rubtsov, and K. Held, Diagrammatic routes to nonlocal correlations beyond dynamical mean field theory, Rev. Mod. Phys. 90, 025003 (2018).
- Poncé et al. [2021] S. Poncé, F. Macheda, E. R. Margine, N. Marzari, N. Bonini, and F. Giustino, First-principles predictions of Hall and drift mobilities in semiconductors, Phys. Rev. Res. 3, 043022 (2021).
- Lihm and Poncé [2024] J.-M. Lihm and S. Poncé, Non-perturbative self-consistent electron-phonon spectral functions and transport [link to be added], Materials Cloud Archive (2024).
- Hamann [2013] D. R. Hamann, Optimized norm-conserving Vanderbilt pseudopotentials, Phys. Rev. B 88, 085117 (2013).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- van Setten et al. [2018] M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, The PseudoDojo: Training and grading a 85 element optimized norm-conserving pseudopotential table, Comput. Phys. Commun. 226, 39 (2018).
- Sohier et al. [2017] T. Sohier, M. Gibertini, M. Calandra, F. Mauri, and N. Marzari, Breakdown of optical phonons’ splitting in two-dimensional materials, Nano Lett. 17, 3758 (2017).
- Giustino et al. [2007] F. Giustino, M. L. Cohen, and S. G. Louie, Electron-phonon interaction using Wannier functions, Phys. Rev. B 76, 165108 (2007).
- Marzari and Vanderbilt [1997] N. Marzari and D. Vanderbilt, Maximally-localized generalized Wannier functions for composite energy bands, Phys. Rev. B 56, 12847 (1997).
- Verdi and Giustino [2015] C. Verdi and F. Giustino, Fröhlich electron-phonon vertex from first principles, Phys. Rev. Lett. 115, 176401 (2015).
- Sjakste et al. [2015] J. Sjakste, N. Vast, M. Calandra, and F. Mauri, Wannier interpolation of the electron-phonon matrix elements in polar semiconductors: Polar-optical coupling in gaas, Phys. Rev. B 92, 054307 (2015).
- Brunin et al. [2020a] G. Brunin, H. P. C. Miranda, M. Giantomassi, M. Royo, M. Stengel, M. J. Verstraete, X. Gonze, G.-M. Rignanese, and G. Hautier, Electron-phonon beyond Fröhlich: Dynamical quadrupoles in polar and covalent solids, Phys. Rev. Lett. 125, 136601 (2020a).
- Brunin et al. [2020b] G. Brunin, H. P. C. Miranda, M. Giantomassi, M. Royo, M. Stengel, M. J. Verstraete, X. Gonze, G.-M. Rignanese, and G. Hautier, Phonon-limited electron mobility in Si, GaAs, and GaP with exact treatment of dynamical quadrupoles, Phys. Rev. B 102, 094308 (2020b).
- Jhalani et al. [2020] V. A. Jhalani, J.-J. Zhou, J. Park, C. E. Dreyer, and M. Bernardi, Piezoelectric electron-phonon interaction from ab initio dynamical quadrupoles: Impact on charge transport in wurtzite GaN, Phys. Rev. Lett. 125, 136602 (2020).
- Park et al. [2020] J. Park, J.-J. Zhou, V. A. Jhalani, C. E. Dreyer, and M. Bernardi, Long-range quadrupole electron-phonon interaction from first principles, Phys. Rev. B 102, 125203 (2020).
- Poncé et al. [2023b] S. Poncé, M. Royo, M. Gibertini, N. Marzari, and M. Stengel, Accurate prediction of Hall mobilities in two-dimensional materials through gauge-covariant quadrupolar contributions, Phys. Rev. Lett. 130, 166301 (2023b).
- Royo and Stengel [2021] M. Royo and M. Stengel, Exact long-range dielectric screening and interatomic force constants in quasi-two-dimensional crystals, Phys. Rev. X 11, 041027 (2021).
- Royo and Stengel [2019] M. Royo and M. Stengel, First-principles theory of spatial dispersion: Dynamical quadrupoles and flexoelectricity, Phys. Rev. X 9, 021050 (2019).
- Gonze et al. [2020] X. Gonze, B. Amadon, G. Antonius, F. Arnardi, L. Baguet, J.-M. Beuken, J. Bieder, F. Bottin, J. Bouchet, E. Bousquet, N. Brouwer, F. Bruneval, G. Brunin, T. Cavignac, J.-B. Charraud, W. Chen, M. Côté, S. Cottenier, J. Denier, G. Geneste, P. Ghosez, M. Giantomassi, Y. Gillet, O. Gingras, D. R. Hamann, G. Hautier, X. He, N. Helbig, N. Holzwarth, Y. Jia, F. Jollet, W. Lafargue-Dit-Hauret, K. Lejaeghere, M. A. Marques, A. Martin, C. Martins, H. P. Miranda, F. Naccarato, K. Persson, G. Petretto, V. Planes, Y. Pouillon, S. Prokhorenko, F. Ricci, G.-M. Rignanese, A. H. Romero, M. M. Schmitt, M. Torrent, M. J. van Setten, B. Van Troeye, M. J. Verstraete, G. Zérah, and J. W. Zwanziger, The Abinit project: Impact, environment and recent developments, Comput. Phys. Commun. 248, 107042 (2020).
- Bezanson et al. [2017] J. Bezanson, A. Edelman, S. Karpinski, and V. B. Shah, Julia: A fresh approach to numerical computing, SIAM Rev. 59, 65 (2017).