### v3

# Nonperturbative self-consistent electron-phonon spectral functions and transport

Abstract
Electron-phonon coupling often dominates the electron spectral functions and carrier transport properties. However, studies of this effect in real materials have largely relied on perturbative one-shot methods due to the lack of a first-principles theoretical and computational framework. Here, we present a self-consistent theory and implementation for the nonperturbative calculations of spectral functions and conductivity due to electron-phonon coupling. Applying this method to monolayer InSe, we demonstrate that self-consistency qualitatively affects the spectral function and transport properties compared to state-of-the-art one-shot calculations and allow one to reconcile calculations with angle-resolved photoemission experiments. The developed method can be widely applied to materials with dominant electron-phonon coupling at moderate computational cost.
Spectral functions are essential in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy (ARPES) [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon (-ph) interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
Ab initio calculation of the -ph self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17], solving the Dyson equation [10] or applying the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25, 26] for spectral functions. However, these approaches often fail, causing divergent quasiparticle dispersion [27, 28, 26] or unphysical spectral properties [26]. Notably, this problem is not limited to strong e-ph coupling: the negative curvature in the quasiparticle dispersion of the Fröhlich model [27] and the divergent linewidth of piezoelectric materials [29] illustrate failures at all coupling strengths. Advanced methods incorporating higher-order -ph interactions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] are limited to model Hamiltonians. Nonperturbative methods using molecular dynamics [45, 46] or special-displacement supercells [47, 48] are difficult to converge [46, 49], and cannot capture nonadiabatic -ph coupling, crucial in polar materials [28, 50].
The electron spectral function also governs charge transport. Within the Green–Kubo theory [51, 52], neglecting current vertex corrections allows direct computation of conductivity from spectral functions [53, 54]. This approach, known as the bubble approximation, has the advantage of including broadening and satellites beyond the quasiparticle approximation [24, 55, 56, 57, 46].
A material of particular interest for nonperturbative -ph coupling is monolayer InSe. While known for high electron mobility [58, 59], its valence band features a flat, Mexican-hat-like dispersion [60, 61, 62, 63, 64, 65, 66], leading to strong -ph interaction. In particular, state-of-the-art one-shot calculations predicted a dynamical splitting of the valence band into two peaks in the spectral function with a gap larger than 150 meV [67]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [68, 69, 70] and the bubble approximation with the one-shot spectral function [67]. Yet, given the strong renormalization, nonperturbative corrections are expected to be essential in monolayer InSe.
In this Letter, we present a nonperturbative ab initio calculation of the -ph spectral function and conductivity. Our approach, the self-consistent (sc) method [36, 37, 38, 39, 40, 41, 42, 43, 44], extends the FM self-energy to include the self-consistently renormalized spectral functions, analogously to the sc method for electron-electron interactions [71]. Using self-consistent spectral functions, we compute conductivity using the bubble approximation. We find that self-consistency significantly affects the spectral function and conductivity of monolayer InSe. This effect is present in any materials and can be explored with the sc method.
The most common approximation for the -ph self-energy is the FM formula [2]:
where is the Brillouin zone volume, the -ph matrix element, and the electron and phonon energies, is a positive infinitesimal broadening, the Fermi–Dirac distribution, , and the Bose-Einstein distribution. We neglect the band off-diagonal self-energy, which is a valid approximation if the bands do not hybridize [74], as in InSe. We refer to Eq. (1) as the -FM self-energy, corresponding to a one-loop diagram (see Fig. LABEL:supp-fig:feynman(a) of Supplemental Material (SM) [72]) with bare electron and phonon Green’s functions, and , as in the approximation [75]. In this Letter, we assume that is obtained from Kohn–Sham density functional theory and that and -ph coupling come from adiabatic density functional perturbation theory [76, 77, 2]. We neglect the renormalization of the phonon Green’s function since we focus on the intrinsic and low doping regime where the phonon renormalization is small [2]. The total self-energy also contains the Debye–Waller (DW) term [5, 4, 3],
In this work, we compute the DW term within the rigid-ion approximation [78].
Using the self-energy, the renormalized electron dispersion can be computed within Rayleigh–Schrödinger (RS) perturbation theory [79, 80, 28, 2, 23] as
The energy-dependent spectral function is computed by solving the Dyson equation
with the chemical potential.
To capture higher-order -ph coupling, one often uses the cumulant approximation. It replaces the Dyson equation with the expansion
where is the Heaviside function and the cumulant function. We adopt the retarded cumulant method [81, 82, 24] as we deal with finite temperatures. The cumulant function is approximated via second-order perturbation theory [81, 21, 82],
Unlike the approximation, which only produces a single satellite [15], the cumulant approach can generate multiple satellite peaks, as seen in experiments [11, 12, 13, 14]. However, it has been shown that the cumulant approximation can lead to unphysically large renormalizations for points away from the band edge [26, 44], similar to the RS approximation [27].
To incorporate nonperturbative -ph effects, we adopt the sc method, where the self-energy is obtained by replacing the bare electron Green’s function in the approximation with the dressed one [see Fig. LABEL:supp-fig:feynman(b) in SM [72]]. The imaginary part of the sc self-energy is given by [56]
The real part is obtained using the Kramers–Kronig relation, with the addition of the real-valued DW term:
where denotes the principal value. The spectral function in Eq. (8) is obtained via the Dyson equation:
Solving Eqs. (8)-(11) self-consistently yields the sc self-energy and spectral function. We also iteratively adjust the temperature-dependent chemical potential to match the total electron occupation.
We remark that it is advantageous to first compute the imaginary part and then use the Kramers–Kronig relation to obtain the real part. The relation between the real parts of and involves a -dependent energy integral [56]. Separating Eq. (8) and Eq. (9) enables performing the integration for the imaginary part first, followed by a -independent energy integral.
In practical ab initio calculations, several hundred bands per atom are needed to converge the real part of the FM self-energy [49]. Calculating this via the Kramers–Kronig transformation is impractical, as it demands evaluating up to eV. We simplify the calculation by treating the contribution of high-energy bands as a static correction. We thus self-consistently compute only for the states inside the active space window , and add the static correction as
Here, is the one-shot FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue . This term can be efficiently computed by solving the linear Sternheimer equation [79, 80] and interpolating using Wannier function perturbation theory (WFPT) [83]. The static corrections in the second line of Eq. (12) are always included during the self-consistent calculation of the sc self-energy, i.e., Eq. (12) replaces Eq. (9). (See Sec. LABEL:supp-sec:active_space of SM [72] for additional details.)
In practice, we initialize the sc self-energy as a constant with meV. We do not use any artificial broadening during the sc iteration, and the choice of does not affect the converged sc self-energy. We iterate until the maximum change of is below meV. The sc method is implemented as an extension of our ElectronPhonon.jl package [84], which uses quantities computed from Quantum ESPRESSO [85], Wannier90 [86], and EPW [87, 88]. Computational details are provided in Sec. LABEL:supp-sec:comp_details of SM [72].
The sc method, also known as the self-consistent Migdal, self-consistent first-Born, or the noncrossing approximation, has been applied to the Holstein [89, 90, 36, 37, 38, 40, 41, 42, 43, 44] and Anderson–Holstein models [39]. This nonperturbative method sums the infinite series of noncrossing rainbow diagrams [see Fig. LABEL:supp-fig:feynman(c) of SM [72]], including those with an arbitrary number of -ph vertices. It is expected to be valid from weak to intermediate coupling but will fail in the strong coupling regime, as in the case of self-trapped small polarons [91]. State-of-the-art ab initio calculations [56, 57] used a one-shot calculation of Eqs. (8, 9) (without ) with the spectral function obtained from dynamical mean-field theory to compute the -ph self-energy. Here, we self-consistently solve Eqs. (8-12) to include nonperturbative -ph effects on the spectral function.
We apply the sc method to monolayer InSe, which has a Mexican-hat-like valance band [60, 61, 62, 64]. This dispersion leads to strong -ph effects in hole-doped InSe monolayers: the one-shot spectral functions show a dynamic splitting of 150 meV or larger [67]. We study whether this large renormalization persists under a self-consistent update of the spectral function.
Figures 1(a)-(d) show the valence band spectral functions of monolayer InSe at 100 K, compared to the ARPES photointensity [73] measured at the same temperature. The RS dispersion has several kinks with inverted curvature near the valence band maximum (VBM), an artifact also observed for the Fröhlich model [27]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot method yields two dynamically split bands with a splitting of 170 meV, consistent with Ref. [67]. However, this large splitting vanishes under self-consistency. The sc spectral function shows a broadened quasiparticle peak with multiple satellites, with a distance between the peaks of 30–50 meV, comparable to the longitudinal optical (LO) phonon energy of 27 meV. An analysis of the difference between and sc self-energies is provided in the End Matter (Fig. 4). Our sc result is consistent with the absence of a 150 meV scale gap in the ARPES photointensity [64, 73]. Although both the sc and cumulant spectral functions are consistent with the experimental observation within the experimental resolution, higher-resolution ARPES may observe the satellites predicted by our sc calculation.
Beyond ARPES predictions that can sometimes be challenging to verify experimentally, the electron spectral function can be directly used to compute carrier conductivity, a simpler quantity to measure. Neglecting the current vertex corrections, we calculate the conductivity in the bubble approximation, which is given by [53, 54]
The corresponding mobility is with the carrier concentration. Equation (13) has been used to calculate the conductivity due to electron-electron interaction [54], -ph interaction [24, 55, 46], or both [56, 57]. Here, we investigate how the self-consistent renormalization of the spectral function affects the dc and ac mobility of monolayer InSe.
To study transport in the intrinsic limit, we set a small carrier concentration, placing the chemical potential within the band gap. In this nondegenerate regime, mobility should be independent of carrier concentration [92, 93]. However, as detailed in the End Matter, we find that both the and cumulant spectral functions yield strongly -dependent mobility (Fig. 5). We attribute this artifact to the slow decay of the spectral function above the band edge, which induces a large number of carriers on the tail of the spectral function. To address this problem, we developed a truncation scheme that removes the tail above a cutoff energy (see Sec. LABEL:supp-sec:decayof SM [72]), which results in a -independent mobility. However, this scheme cannot be applied to the cumulant spectral function of InSe due to its large broadening. The sc method does not suffer from this problem and yields an -independent mobility.
Figure 2(a) shows the intrinsic dc mobility of InSe. While the BTE and self-energy relaxation time approximation (SERTA) mobilities [94] are nearly identical and agree with the results from Refs. [69, 70], the bubble mobility with the sc spectral function is 3 to 4 times larger. This suggests that the quasiparticle approximation significantly underestimates the mobility. Using the one-shot spectral function instead of the sc one leads to an overestimation by more than an order of magnitude, and the mobility strongly depends on the artificial broadening , varying inversely with [see Fig. LABEL:supp-fig:dc_sc_broadening of SM [72]]. These results can be understood in terms of the effective electron scattering rates shown in Fig. 2(b), which are inversely proportional to the mobility. The SERTA and BTE mobilities are linked to the the imaginary part of the self-energy at the bare band energy,
For the spectral functions, we define the effective scattering rates using the self-energy at the renormalized quasiparticle energy,
The renormalized quasiparticle energies, and , are the energies where the spectral functions and are maximal. Near the VBM, the BTE scattering rates exhibit a peak, while the sc result shows a smoother variation and smaller values due to the self-consistent broadening. For the one-shot self-energy, the scattering rates near the VBM are tiny because the renormalized valence bands lie within the bare band gap where the imaginary part of the self-energy is small. This behavior is an artifact of the one-shot approximation.
Finally, we show in Fig. 3 the ac mobility calculated using the and sc spectral functions. For the one-shot spectral function, we observe a peak around 150–250 meV, corresponding to the gap shown in Fig. 1(c), as in Ref. [67]. This feature persists with a larger artificial broadening (see Fig. LABEL:supp-fig:ac_sc_broadening of SM [72]). However, with the self-consistent spectral function, the unphysical peaks are smoothed out due to the reduced splitting and the broader Drude peak.
The self-consistent framework developed in this work offers a versatile tool for analyzing the nonperturbative effect of -ph coupling with moderate computational effort, bridging the gap between one-shot approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, this framework can be extended to other properties such as superconductivity and indirect absorption [95, 96]. Our work also lays the groundwork for applying diagrammatic many-body approaches to -ph systems. The sc method could be extended to explore contributions beyond the Fan–Migdal diagram. Notably, the development of a fully self-consistent scheme using the renormalized phonon Green’s function [36, 37, 97], the inclusion of the dynamical screening [98, 99, 100], and the consideration of -ph vertex correction [40, 42, 101] is a promising avenue for future work. Combining the sc formalism with dynamical mean-field theory [102, 103] and its diagrammatic extensions [104] to study transport from the interplay of electron-electron [54] and -ph coupling [105] beyond the one-shot approximation [56, 57] is another interesting direction for future research.
Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73], and the anonymous referees who provided invaluable insights that led to an improved version of this work. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247. Data availability—The data that support the findings of this article are openly available on the Materials Cloud Archive [106].References
- Damascelli et al. [2003] A. Damascelli, Z. Hussain, and Z.-X. Shen, Angle-resolved photoemission studies of the cuprate superconductors, Rev. Mod. Phys. 75, 473 (2003).
- Giustino [2017] F. Giustino, Electron-phonon interactions from first principles, Rev. Mod. Phys. 89, 015003 (2017).
- Allen and Cardona [1983] P. B. Allen and M. Cardona, Temperature dependence of the direct gap of Si and Ge, Phys. Rev. B 27, 4760 (1983).
- Allen and Cardona [1981] P. B. Allen and M. Cardona, Theory of the temperature dependence of the direct gap of germanium, Phys. Rev. B 23, 1495 (1981).
- Allen and Heine [1976] P. B. Allen and V. Heine, Theory of the temperature dependence of electronic band structures, J. Phys. Condens. Matter 9, 2305 (1976).
- Marini [2008] A. Marini, Ab initio finite-temperature excitons, Phys. Rev. Lett. 101, 106405 (2008).
- Giustino et al. [2010] F. Giustino, S. G. Louie, and M. L. Cohen, Electron-phonon renormalization of the direct band gap of diamond, Phys. Rev. Lett. 105, 265501 (2010).
- Valla et al. [1999] T. Valla, A. V. Fedorov, P. D. Johnson, and S. L. Hulbert, Many-body effects in angle-resolved photoemission: Quasiparticle energy and lifetime of a Mo(110) surface state, Phys. Rev. Lett. 83, 2085 (1999).
- Eiguren et al. [2003] A. Eiguren, S. de Gironcoli, E. V. Chulkov, P. M. Echenique, and E. Tosatti, Electron-phonon interaction at the Be(0001) surface, Phys. Rev. Lett. 91, 166803 (2003).
- Engelsberg and Schrieffer [1963] S. Engelsberg and J. R. Schrieffer, Coupled electron-phonon system, Phys. Rev. 131, 993 (1963).
- Moser et al. [2013] S. Moser, L. Moreschini, J. Jaćimović, O. S. Barišić, H. Berger, A. Magrez, Y. J. Chang, K. S. Kim, A. Bostwick, E. Rotenberg, L. Forró, and M. Grioni, Tunable polaronic conduction in anatase , Phys. Rev. Lett. 110, 196403 (2013).
- Chen et al. [2015] C. Chen, J. Avila, E. Frantzeskakis, A. Levy, and M. C. Asensio, Observation of a two-dimensional liquid of Fröhlich polarons at the bare SrTiO3 surface, Nat. Commun. 6, 8585 (2015).
- Cancellieri et al. [2016] C. Cancellieri, A. S. Mishchenko, U. Aschauer, A. Filippetti, C. Faber, O. Barišić, V. Rogalev, T. Schmitt, N. Nagaosa, and V. N. Strocov, Polaronic metal state at the LaAlO3/SrTiO3 interface, Nat. Commun. 7, 10386 (2016).
- Wang et al. [2016] Z. Wang, S. McKeown Walker, A. Tamai, Y. Wang, Z. Ristic, F. Y. Bruno, A. De La Torre, S. Riccò, N. C. Plumb, M. Shi, P. Hlawenka, J. Sánchez-Barriga, A. Varykhalov, T. K. Kim, M. Hoesch, P. D. C. King, W. Meevasana, U. Diebold, J. Mesot, B. Moritz, T. P. Devereaux, M. Radovic, and F. Baumberger, Tailoring the nature and strength of electron-phonon interactions in the SrTiO3(001) 2D electron liquid, Nat. Mater. 15, 835 (2016).
- Verdi et al. [2017] C. Verdi, F. Caruso, and F. Giustino, Origin of the crossover from polarons to Fermi liquids in transition metal oxides, Nat. Commun. 8, 15769 (2017).
- Migdal [1958] A. Migdal, Interaction between electrons and lattice vibrations in a normal metal, Sov. Phys. JETP 7, 996 (1958).
- Fan [1951] H. Y. Fan, Temperature dependence of the energy gap in semiconductors, Phys. Rev. 82, 900 (1951).
- Gunnarsson et al. [1994] O. Gunnarsson, V. Meden, and K. Schönhammer, Corrections to Migdal’s theorem for spectral functions: A cumulant treatment of the time-dependent Green’s function, Phys. Rev. B 50, 10462 (1994).
- Aryasetiawan et al. [1996] F. Aryasetiawan, L. Hedin, and K. Karlsson, Multiple plasmon satellites in Na and Al spectral functions from ab initio cumulant expansion, Phys. Rev. Lett. 77, 2268 (1996).
- Guzzo et al. [2011] M. Guzzo, G. Lani, F. Sottile, P. Romaniello, M. Gatti, J. J. Kas, J. J. Rehr, M. G. Silly, F. Sirotti, and L. Reining, Valence electron photoemission spectrum of semiconductors: Ab initio description of multiple satellites, Phys. Rev. Lett. 107, 166401 (2011).
- Story et al. [2014] S. M. Story, J. J. Kas, F. D. Vila, M. J. Verstraete, and J. J. Rehr, Cumulant expansion for phonon contributions to the electron spectral function, Phys. Rev. B 90, 195135 (2014).
- Gumhalter et al. [2016] B. Gumhalter, V. Kovač, F. Caruso, H. Lambert, and F. Giustino, On the combined use of approximation and cumulant expansion in the calculations of quasiparticle spectra: The paradigm of Si valence bands, Phys. Rev. B 94, 035103 (2016).
- Nery et al. [2018] J. P. Nery, P. B. Allen, G. Antonius, L. Reining, A. Miglio, and X. Gonze, Quasiparticles and phonon satellites in spectral functions of semiconductors and insulators: Cumulants applied to the full first-principles theory and the Fröhlich polaron, Phys. Rev. B 97, 115145 (2018).
- Zhou and Bernardi [2019] J.-J. Zhou and M. Bernardi, Predicting charge transport in the presence of polarons: The beyond-quasiparticle regime in SrTiO3, Phys. Rev. Res. 1, 033138 (2019).
- de Abreu et al. [2022] J. C. de Abreu, J. P. Nery, M. Giantomassi, X. Gonze, and M. J. Verstraete, Spectroscopic signatures of nonpolarons: the case of diamond, Phys. Chem. Chem. Phys. 24, 12580 (2022).
- Kandolf et al. [2022] N. Kandolf, C. Verdi, and F. Giustino, Many-body Green’s function approaches to the doped Fröhlich solid: Exact solutions and anomalous mass enhancement, Phys. Rev. B 105, 085148 (2022).
- Larsen [1966] D. M. Larsen, Polaron energy spectrum, Phys. Rev. 144, 697 (1966).
- Poncé et al. [2015] S. Poncé, Y. Gillet, J. Laflamme Janssen, A. Marini, M. Verstraete, and X. Gonze, Temperature dependence of the electronic structure of semiconductors and insulators, J. Chem. Phys. 143, 102813 (2015).
- Lihm et al. [2024] J.-M. Lihm, S. Poncé, and C.-H. Park, Self-consistent electron lifetimes for electron-phonon scattering, Phys. Rev. B 110, L121106 (2024).
- Prokof’ev and Svistunov [1998] N. V. Prokof’ev and B. V. Svistunov, Polaron problem by diagrammatic quantum Monte Carlo, Phys. Rev. Lett. 81, 2514 (1998).
- Jeckelmann and White [1998] E. Jeckelmann and S. R. White, Density-matrix renormalization-group study of the polaron problem in the Holstein model, Phys. Rev. B 57, 6376 (1998).
- Berciu [2006] M. Berciu, Green’s function of a dressed particle, Phys. Rev. Lett. 97, 036402 (2006).
- Bonča et al. [2019] J. Bonča, S. A. Trugman, and M. Berciu, Spectral function of the Holstein polaron at finite temperature, Phys. Rev. B 100, 094307 (2019).
- Carbone et al. [2021] M. R. Carbone, D. R. Reichman, and J. Sous, Numerically exact generalized Green’s function cluster expansions for electron-phonon problems, Phys. Rev. B 104, 035106 (2021).
- Mahajan et al. [2024] A. Mahajan, P. J. Robinson, J. Lee, and D. R. Reichman, Structure and dynamics of electron-phonon coupled systems using neural quantum states, arXiv:2405.08701 (2024).
- Säkkinen et al. [2015a] N. Säkkinen, Y. Peng, H. Appel, and R. Van Leeuwen, Many-body Green’s function theory for electron-phonon interactions: Ground state properties of the Holstein dimer, J. Chem. Phys. 143, 234101 (2015a).
- Säkkinen et al. [2015b] N. Säkkinen, Y. Peng, H. Appel, and R. Van Leeuwen, Many-body Green’s function theory for electron-phonon interactions: The Kadanoff-Baym approach to spectral properties of the Holstein dimer, J. Chem. Phys. 143, 234102 (2015b).
- Rademaker et al. [2016] L. Rademaker, Y. Wang, T. Berlijn, and S. Johnston, Enhanced superconductivity due to forward scattering in fese thin films on SrTiO3 substrates, New J. Phys. 18, 022001 (2016).
- Chen et al. [2016] H.-T. Chen, G. Cohen, A. J. Millis, and D. R. Reichman, Anderson-Holstein model in two flavors of the noncrossing approximation, Phys. Rev. B 93, 174309 (2016).
- Esterlis et al. [2018] I. Esterlis, B. Nosarzewski, E. W. Huang, B. Moritz, T. P. Devereaux, D. J. Scalapino, and S. A. Kivelson, Breakdown of the Migdal-Eliashberg theory: A determinant quantum Monte Carlo study, Phys. Rev. B 97, 140501 (2018).
- Dee et al. [2019] P. M. Dee, K. Nakatsukasa, Y. Wang, and S. Johnston, Temperature-filling phase diagram of the two-dimensional Holstein model in the thermodynamic limit by self-consistent Migdal approximation, Phys. Rev. B 99, 024514 (2019).
- Dee et al. [2020] P. M. Dee, J. Coulter, K. G. Kleiner, and S. Johnston, Relative importance of nonlinear electron-phonon coupling and vertex corrections in the Holstein model, Commun. Phys. 3, 145 (2020).
- Mitrić et al. [2022] P. Mitrić, V. Janković, N. Vukmirović, and D. Tanasković, Spectral functions of the Holstein polaron: Exact and approximate solutions, Phys. Rev. Lett. 129, 096401 (2022).
- Mitrić et al. [2023] P. Mitrić, V. Janković, N. Vukmirović, and D. Tanasković, Cumulant expansion in the Holstein model: Spectral functions and mobility, Phys. Rev. B 107, 125165 (2023).
- Zacharias et al. [2020] M. Zacharias, M. Scheffler, and C. Carbogno, Fully anharmonic nonperturbative theory of vibronically renormalized electronic band structures, Phys. Rev. B 102, 045126 (2020).
- Quan et al. [2024] J. Quan, C. Carbogno, and M. Scheffler, Carrier mobility of strongly anharmonic materials from first principles, Phys. Rev. B 110, 235202 (2024).
- Zacharias and Giustino [2016] M. Zacharias and F. Giustino, One-shot calculation of temperature-dependent optical spectra and phonon-induced band-gap renormalization, Phys. Rev. B 94, 075125 (2016).
- Zacharias and Giustino [2020] M. Zacharias and F. Giustino, Theory of the special displacement method for electronic structure calculations at finite temperature, Phys. Rev. Res. 2, 013357 (2020).
- Poncé et al. [2024] S. Poncé, J.-M. Lihm, and C.-H. Park, Zero-point renormalization of the bandgap, mass enhancement, and spectral functions: Validation of methods and verification of first-principles codes, arXiv:2410.14319 (2024).
- Miglio et al. [2020] A. Miglio, V. Brousseau-Couture, E. Godbout, G. Antonius, Y.-H. Chan, S. G. Louie, M. Côté, M. Giantomassi, and X. Gonze, Predominance of non-adiabatic effects in zero-point renormalization of the electronic band gap, npj Comput. Mater. 6, 167 (2020).
- Green [1954] M. S. Green, Markoff random processes and the statistical mechanics of time-dependent phenomena. II. irreversible processes in fluids, J. Chem. Phys. 22, 398 (1954).
- Kubo [1957] R. Kubo, Statistical-mechanical theory of irreversible processes. I. general theory and simple applications to magnetic and conduction problems, J. Phys. Soc. Jpn. 12, 570 (1957).
- Mahan [2000] G. D. Mahan, Many-particle physics, 3rd ed. (Kluwer Academic / Plenum Publishers, New York, 2000).
- Basov et al. [2011] D. N. Basov, R. D. Averitt, D. van der Marel, M. Dressel, and K. Haule, Electrodynamics of correlated electron materials, Rev. Mod. Phys. 83, 471 (2011).
- Chang et al. [2022] B. K. Chang, J.-J. Zhou, N.-E. Lee, and M. Bernardi, Intermediate polaronic charge transport in organic crystals from a many-body first-principles approach, npj Comput. Mater. 8, 63 (2022).
- Abramovitch et al. [2023] D. J. Abramovitch, J.-J. Zhou, J. Mravlje, A. Georges, and M. Bernardi, Combining electron-phonon and dynamical mean-field theory calculations of correlated materials: Transport in the correlated metal , Phys. Rev. Mater. 7, 093801 (2023).
- Abramovitch et al. [2024] D. J. Abramovitch, J. Mravlje, J.-J. Zhou, A. Georges, and M. Bernardi, Respective roles of electron-phonon and electron-electron interactions in the transport and quasiparticle properties of , Phys. Rev. Lett. 133, 186501 (2024).
- Bandurin et al. [2017] D. A. Bandurin, A. V. Tyurnina, G. L. Yu, A. Mishchenko, V. Zólyomi, S. V. Morozov, R. K. Kumar, R. V. Gorbachev, Z. R. Kudrynskyi, S. Pezzini, et al., High electron mobility, quantum Hall effect and anomalous optical response in atomically thin InSe, Nat. Nanotechnol. 12, 223 (2017).
- Wan et al. [2023] W. Wan, R. Guo, Y. Ge, and Y. Liu, Carrier and phonon transport in 2d InSe and its Janus structures, J. Condens. Matter Phys. 35, 133001 (2023).
- Zólyomi et al. [2014] V. Zólyomi, N. D. Drummond, and V. I. Fal’ko, Electrons and phonons in single layers of hexagonal indium chalcogenides from ab initio calculations, Phys. Rev. B 89, 205416 (2014).
- Rybkovskiy et al. [2014] D. V. Rybkovskiy, A. V. Osadchy, and E. D. Obraztsova, Transition from parabolic to ring-shaped valence band maximum in few-layer GaS, GaSe, and InSe, Phys. Rev. B 90, 235302 (2014).
- Debbichi et al. [2015] L. Debbichi, O. Eriksson, and S. Lebégue, Two-dimensional indium selenides compounds: An ab initio study, J. Phys. Chem. Lett. 6, 3098 (2015).
- Magorrian et al. [2016] S. J. Magorrian, V. Zólyomi, and V. I. Fal’ko, Electronic and optical properties of two-dimensional InSe from a DFT-parametrized tight-binding model, Phys. Rev. B 94, 245431 (2016).
- Kibirev et al. [2018] I. A. Kibirev, A. V. Matetskiy, A. V. Zotov, and A. A. Saranin, Thickness-dependent transition of the valence band shape from parabolic to Mexican-hat-like in the MBE grown InSe ultrathin films, Appl. Phys. Lett. 112, 191602 (2018).
- Graham et al. [2020] A. J. Graham, J. Zultak, M. J. Hamer, V. Zolyomi, S. Magorrian, A. Barinov, V. Kandyba, A. Giampietri, A. Locatelli, F. Genuzio, N. C. Teutsch, C. Salazar, N. D. M. Hine, V. I. Fal’ko, R. V. Gorbachev, and N. R. Wilson, Ghost anti-crossings caused by interlayer umklapp hybridization of bands in 2D heterostructures, 2D Materials 8, 015016 (2020).
- Stepanov et al. [2022] E. A. Stepanov, V. Harkov, M. Rösner, A. I. Lichtenstein, M. I. Katsnelson, and A. N. Rudenko, Coexisting charge density wave and ferromagnetic instabilities in monolayer InSe, npj Comput. Mater. 8, 118 (2022).
- Lugovskoi et al. [2019] A. V. Lugovskoi, M. I. Katsnelson, and A. N. Rudenko, Strong electron-phonon coupling and its influence on the transport and optical properties of hole-doped single-layer InSe, Phys. Rev. Lett. 123, 176401 (2019).
- Das et al. [2019] P. Das, D. Wickramaratne, B. Debnath, G. Yin, and R. K. Lake, Charged impurity scattering in two-dimensional materials with ring-shaped valence bands: GaS, GaSe, InS, and InSe, Phys. Rev. B 99, 085409 (2019).
- Li et al. [2019] W. Li, S. Poncé, and F. Giustino, Dimensional crossover in the carrier mobility of two-dimensional semiconductors: The case of InSe, Nano Lett. 19, 1774 (2019).
- Poncé et al. [2023a] S. Poncé, M. Royo, M. Stengel, N. Marzari, and M. Gibertini, Long-range electrostatic contribution to electron-phonon couplings and mobilities of two-dimensional and bulk materials, Phys. Rev. B 107, 155424 (2023a).
- Shishkin and Kresse [2007] M. Shishkin and G. Kresse, Self-consistent calculations for semiconductors and insulators, Phys. Rev. B 75, 235102 (2007).
- Note [1] See Supplemental Material [URL will be inserted by publisher], which includes Refs. [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125], for the derivation of the formula, computational details, and additional results.
- Hamer et al. [2019] M. J. Hamer, J. Zultak, A. V. Tyurnina, V. Zólyomi, D. Terry, A. Barinov, A. Garner, J. Donoghue, A. P. Rooney, V. Kandyba, A. Giampietri, A. Graham, N. Teutsch, X. Xia, M. Koperski, S. J. Haigh, V. I. Fal’ko, R. V. Gorbachev, and N. R. Wilson, Indirect to direct gap crossover in two-dimensional inse revealed by angle-resolved photoemission spectroscopy, ACS Nano 13, 2136 (2019).
- Lihm and Park [2020] J.-M. Lihm and C.-H. Park, Phonon-induced renormalization of electron wave functions, Phys. Rev. B 101, 121102 (2020).
- Hybertsen and Louie [1986] M. S. Hybertsen and S. G. Louie, Electron correlation in semiconductors and insulators: Band gaps and quasiparticle energies, Phys. Rev. B 34, 5390 (1986).
- Gonze and Lee [1997] X. Gonze and C. Lee, Dynamical matrices, born effective charges, dielectric permittivity tensors, and interatomic force constants from density-functional perturbation theory, Phys. Rev. B 55, 10355 (1997).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Phonons and related crystal properties from density-functional perturbation theory, Rev. Mod. Phys. 73, 515 (2001).
- Poncé et al. [2014] S. Poncé, G. Antonius, Y. Gillet, P. Boulanger, J. Laflamme Janssen, A. Marini, M. Côté, and X. Gonze, Temperature dependence of electronic eigenenergies in the adiabatic harmonic approximation, Phys. Rev. B 90, 214304 (2014).
- Gonze et al. [2011] X. Gonze, P. Boulanger, and M. Côté, Theoretical approaches to the temperature and zero-point motion effects on the electronic band structure, Ann. Phys. 523, 168 (2011).
- Poncé et al. [2014] S. Poncé, G. Antonius, P. Boulanger, E. Cannuccia, A. Marini, M. Côté, and X. Gonze, Verification of first-principles codes: Comparison of total energies, phonon frequencies, electron-phonon coupling and zero-point motion correction to the gap between ABINIT and QE/Yambo, Comput. Mater. Sci. 83, 341 (2014).
- Kas et al. [2014] J. J. Kas, J. J. Rehr, and L. Reining, Cumulant expansion of the retarded one-electron Green function, Phys. Rev. B 90, 085112 (2014).
- Kas and Rehr [2017] J. J. Kas and J. J. Rehr, Finite temperature Green’s function approach for excited state and thermodynamic properties of cool to warm dense matter, Phys. Rev. Lett. 119, 176403 (2017).
- Lihm and Park [2021] J.-M. Lihm and C.-H. Park, Wannier function perturbation theory: Localized representation and interpolation of wave function perturbation, Phys. Rev. X 11, 041053 (2021).
- Lihm [2024] J.-M. Lihm, ElectronPhonon.jl (2024).
- Giannozzi et al. [2017] P. Giannozzi et al., Advanced capabilities for materials modelling with Quantum ESPRESSO, J. Condens. Matter Phys. 29, 465901 (2017).
- Pizzi et al. [2020] G. Pizzi et al., Wannier90 as a community code: New features and applications, J. Condens. Matter Phys. 32, 165902 (2020).
- Poncé et al. [2016] S. Poncé, E. Margine, C. Verdi, and F. Giustino, EPW: Electron-phonon coupling, transport and superconducting properties using maximally localized Wannier functions, Comput. Phys. Commun. 209, 116 (2016).
- Lee et al. [2023] H. Lee, S. Poncé, K. Bushick, S. Hajinazar, J. Lafuente-Bartolome, J. Leveillee, C. Lian, J.-M. Lihm, F. Macheda, H. Mori, H. Paudyal, W. H. Sio, S. Tiwari, M. Zacharias, X. Zhang, N. Bonini, E. Kioupakis, E. R. Margine, and F. Giustino, Electron–phonon physics from first principles using the EPW code, npj Comput. Mater. 9, 156 (2023).
- Capone and Ciuchi [2003] M. Capone and S. Ciuchi, Polaron crossover and bipolaronic metal-insulator transition in the half-filled Holstein model, Phys. Rev. Lett. 91, 186405 (2003).
- Bauer et al. [2011] J. Bauer, J. E. Han, and O. Gunnarsson, Quantitative reliability study of the Migdal-Eliashberg theory for strong electron-phonon coupling in superconductors, Phys. Rev. B 84, 184531 (2011).
- Franchini et al. [2021] C. Franchini, M. Reticcioli, M. Setvin, and U. Diebold, Polarons in materials, Nat. Rev. Mater. 10.1038/s41578-021-00289-w (2021).
- Sze and Ng [2006] S. M. Sze and K. K. Ng, Physics of Semiconductor Devices (John Wiley & Sons, Inc., Hoboken, NJ, USA, 2006).
- Lundstrom [2010] M. Lundstrom, Fundamentals of Carrier Transport, 2nd ed. (Cambridge University Press, Cambridge, 2010).
- Poncé et al. [2018] S. Poncé, E. R. Margine, and F. Giustino, Towards predictive many-body calculations of phonon-limited carrier mobilities in semiconductors, Phys. Rev. B 97, 121201 (2018).
- Noffsinger et al. [2012] J. Noffsinger, E. Kioupakis, C. G. Van de Walle, S. G. Louie, and M. L. Cohen, Phonon-assisted optical absorption in silicon from first principles, Phys. Rev. Lett. 108, 167402 (2012).
- Tiwari et al. [2024] S. Tiwari, E. Kioupakis, J. Menendez, and F. Giustino, Unified theory of optical absorption and luminescence including both direct and phonon-assisted processes, Phys. Rev. B 109, 195127 (2024).
- Berges et al. [2023] J. Berges, N. Girotto, T. Wehling, N. Marzari, and S. Poncé, Phonon self-energy corrections: To screen, or not to screen, Phys. Rev. X 13, 041009 (2023).
- Macheda et al. [2024] F. Macheda, F. Mauri, and T. Sohier, Ab initio Van der Waals electrodynamics: Polaritons and electron scattering from plasmons and phonons in BN-capped graphene, Phys. Rev. B 110, 115407 (2024).
- Lihm and Park [2024] J.-M. Lihm and C.-H. Park, Plasmon-phonon hybridization in doped semiconductors from first principles, Phys. Rev. Lett. 133, 116402 (2024).
- Krsnik et al. [2024] J. Krsnik, D. Novko, and O. S. Barišić, Superconductivity in two-dimensional systems enhanced by nonadiabatic phonon-production effects, Phys. Rev. B 110, L180505 (2024).
- Lee et al. [2020] N.-E. Lee, J.-J. Zhou, H.-Y. Chen, and M. Bernardi, Ab initio electron-two-phonon scattering in GaAs from next-to-leading order perturbation theory, Nat. Commun. 11, 1607 (2020).
- Georges et al. [1996] A. Georges, G. Kotliar, W. Krauth, and M. J. Rozenberg, Dynamical mean-field theory of strongly correlated fermion systems and the limit of infinite dimensions, Rev. Mod. Phys. 68, 13 (1996).
- Kotliar et al. [2006] G. Kotliar, S. Y. Savrasov, K. Haule, V. S. Oudovenko, O. Parcollet, and C. A. Marianetti, Electronic structure calculations with dynamical mean-field theory, Rev. Mod. Phys. 78, 865 (2006).
- Rohringer et al. [2018] G. Rohringer, H. Hafermann, A. Toschi, A. A. Katanin, A. E. Antipov, M. I. Katsnelson, A. I. Lichtenstein, A. N. Rubtsov, and K. Held, Diagrammatic routes to nonlocal correlations beyond dynamical mean field theory, Rev. Mod. Phys. 90, 025003 (2018).
- Poncé et al. [2021] S. Poncé, F. Macheda, E. R. Margine, N. Marzari, N. Bonini, and F. Giustino, First-principles predictions of Hall and drift mobilities in semiconductors, Phys. Rev. Res. 3, 043022 (2021).
- Lihm and Poncé [2025] J.-M. Lihm and S. Poncé, Non-perturbative self-consistent electron-phonon spectral functions and transport, Materials Cloud Archive 10.24435/materialscloud:k1-rt (2025).
- Dunn [1975] D. Dunn, Electron-phonon interactions in an insulator, Can. J. Phys. 53, 321 (1975).
- Nery and Mauri [2022] J. P. Nery and F. Mauri, Nonperturbative green’s function method to determine the electronic spectral function due to electron-phonon interactions: Application to a graphene model from weak to strong coupling, Phys. Rev. B 105, 245120 (2022).
- Hamann [2013] D. R. Hamann, Optimized norm-conserving Vanderbilt pseudopotentials, Phys. Rev. B 88, 085117 (2013).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- van Setten et al. [2018] M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, The PseudoDojo: Training and grading a 85 element optimized norm-conserving pseudopotential table, Comput. Phys. Commun. 226, 39 (2018).
- Sohier et al. [2017] T. Sohier, M. Gibertini, M. Calandra, F. Mauri, and N. Marzari, Breakdown of optical phonons’ splitting in two-dimensional materials, Nano Lett. 17, 3758 (2017).
- Royo and Stengel [2019] M. Royo and M. Stengel, First-principles theory of spatial dispersion: Dynamical quadrupoles and flexoelectricity, Phys. Rev. X 9, 021050 (2019).
- Gonze et al. [2020] X. Gonze, B. Amadon, G. Antonius, F. Arnardi, L. Baguet, J.-M. Beuken, J. Bieder, F. Bottin, J. Bouchet, E. Bousquet, N. Brouwer, F. Bruneval, G. Brunin, T. Cavignac, J.-B. Charraud, W. Chen, M. Côté, S. Cottenier, J. Denier, G. Geneste, P. Ghosez, M. Giantomassi, Y. Gillet, O. Gingras, D. R. Hamann, G. Hautier, X. He, N. Helbig, N. Holzwarth, Y. Jia, F. Jollet, W. Lafargue-Dit-Hauret, K. Lejaeghere, M. A. Marques, A. Martin, C. Martins, H. P. Miranda, F. Naccarato, K. Persson, G. Petretto, V. Planes, Y. Pouillon, S. Prokhorenko, F. Ricci, G.-M. Rignanese, A. H. Romero, M. M. Schmitt, M. Torrent, M. J. van Setten, B. Van Troeye, M. J. Verstraete, G. Zérah, and J. W. Zwanziger, The Abinit project: Impact, environment and recent developments, Comput. Phys. Commun. 248, 107042 (2020).
- Giustino et al. [2007] F. Giustino, M. L. Cohen, and S. G. Louie, Electron-phonon interaction using Wannier functions, Phys. Rev. B 76, 165108 (2007).
- Marzari and Vanderbilt [1997] N. Marzari and D. Vanderbilt, Maximally-localized generalized Wannier functions for composite energy bands, Phys. Rev. B 56, 12847 (1997).
- Verdi and Giustino [2015] C. Verdi and F. Giustino, Fröhlich electron-phonon vertex from first principles, Phys. Rev. Lett. 115, 176401 (2015).
- Sjakste et al. [2015] J. Sjakste, N. Vast, M. Calandra, and F. Mauri, Wannier interpolation of the electron-phonon matrix elements in polar semiconductors: Polar-optical coupling in GaAs, Phys. Rev. B 92, 054307 (2015).
- Brunin et al. [2020a] G. Brunin, H. P. C. Miranda, M. Giantomassi, M. Royo, M. Stengel, M. J. Verstraete, X. Gonze, G.-M. Rignanese, and G. Hautier, Electron-phonon beyond Fröhlich: Dynamical quadrupoles in polar and covalent solids, Phys. Rev. Lett. 125, 136601 (2020a).
- Brunin et al. [2020b] G. Brunin, H. P. C. Miranda, M. Giantomassi, M. Royo, M. Stengel, M. J. Verstraete, X. Gonze, G.-M. Rignanese, and G. Hautier, Phonon-limited electron mobility in Si, GaAs, and GaP with exact treatment of dynamical quadrupoles, Phys. Rev. B 102, 094308 (2020b).
- Jhalani et al. [2020] V. A. Jhalani, J.-J. Zhou, J. Park, C. E. Dreyer, and M. Bernardi, Piezoelectric electron-phonon interaction from ab initio dynamical quadrupoles: Impact on charge transport in wurtzite GaN, Phys. Rev. Lett. 125, 136602 (2020).
- Park et al. [2020] J. Park, J.-J. Zhou, V. A. Jhalani, C. E. Dreyer, and M. Bernardi, Long-range quadrupole electron-phonon interaction from first principles, Phys. Rev. B 102, 125203 (2020).
- Poncé et al. [2023b] S. Poncé, M. Royo, M. Gibertini, N. Marzari, and M. Stengel, Accurate prediction of Hall mobilities in two-dimensional materials through gauge-covariant quadrupolar contributions, Phys. Rev. Lett. 130, 166301 (2023b).
- Royo and Stengel [2021] M. Royo and M. Stengel, Exact long-range dielectric screening and interatomic force constants in quasi-two-dimensional crystals, Phys. Rev. X 11, 041027 (2021).
- Bezanson et al. [2017] J. Bezanson, A. Edelman, S. Karpinski, and V. B. Shah, Julia: A fresh approach to numerical computing, SIAM Rev. 59, 65 (2017).
End Matter
Comparison of the self-energies—Figure 4 compares the and sc self-energies at 100 K and 300 K. The one-shot self-energy shows a sharp peak in the imaginary part at , where is the LO phonon frequency. This peak, associated with phonon emission, is significantly damped and smeared in the sc self-energy due to the self-consistent broadening of the spectral function. This self-consistent regularization becomes more pronounced at higher temperatures.
At 300 K, the sc self-energy reaches around 0.1 eV. While this value is smaller than the full bandwidth of the valence band of InSe (1.2 eV; see Fig. LABEL:supp-fig:InSe_itp of SM [72]), it is larger than the width of the Mexican-hat feature at the valence band edge (0.07 eV). Thus, the self-energy leads to a significant renormalization near the valence band edge, qualitatively altering the spectral function.
Tail of the spectral functions—Compared to the and cumulant spectral functions, the sc spectral functions show a clear qualitative difference in insulators and nondegenerate semiconductors: the former exhibits an unphysically long tail in the midgap region, which is absent in the latter. Figures 5(a, b) show the density of states (DOS)
and the hole DOS (). In Fig. 5(a), the DOS from the and the cumulant methods shows a slow decay (gray shaded curves). This is a result of the Lorentzian broadening of the bare Green’s function (see Sec. LABEL:supp-sec:decay of SM [72] for a detailed derivation.) Importantly, reducing the artificial broadening reduces the amplitude of the tail but does not change the scaling. When multiplied by the hole occupation factor, which increases exponentially with energy up to the chemical potential, the hole DOS continues to increase above the renormalized band edge, peaking at the chemical potential [Fig. 5(b)]. As a result, in nondegenerate semiconductors, the and cumulant methods incorrectly predict an unphysically large amount of carriers in the midgap region, where the spectral function is small. In contrast, the sc spectral function decays faster than (red shaded curve), and the hole DOS peaks at the renormalized band edge as expected.
Doping dependence of the conductivity—Figure 5(c) shows the doping dependence of the hole mobility of monolayer InSe at . Surprisingly, the (green plus markers) and cumulant mobility (blue squares) show significant changes as a function of hole concentration. This result contrasts sharply with what is typically observed in nondegenerate semiconductors with dominant -ph scattering [92, 93]. We attribute this artifact to the unphysical tail in the spectral functions, which places most of the holes near the chemical potential, where the actual spectral function is small. Because of this strong doping dependence, the intrinsic (low doping limit) mobility cannot be properly calculated using the and cumulant approximations. To alleviate this problem, we developed a truncation scheme [see Eq. (LABEL:supp-eq:trunc_occ_p_doped) of SM [72]] that removes the tail of the spectral function beyond the local minima in the hole DOS. Then, we obtain a doping-independent Bubble- mobility (green crosses). However, this truncation could not be applied to the cumulant spectral functions of InSe due to its large broadening.
In contrast, the sc mobility (orange circles) as well as the BTE and SERTA mobility (blue and red triangles) remain doping-independent in the nondegenerate regime (white background). Doping dependence is observed only in the metallic regime (gray background), where the chemical potential lies within the valence band. Therefore, it is highly desirable to use the sc method, rather than the or cumulant approximations, for studying transport in nondegenerate semiconductors.